OpenStudy (anonymous):

Help in Algebra 2 having trouble Factoring Let f(x) = 4x - 7 and g(x) = 2x - 3. Find (f*g)(4) a. 9 b. 13 c. 5 d. 15

OpenStudy (haseeb96):

f(x) * g(x)= (f*g)(x) (4x-7)(2x-3) = 8x^2 -12x -14x +21 = 8x^2 -26x +21 now(f*g)(x) = 8x^2 -26x +21 replace x with 4 and get your final answer and show me what u get

OpenStudy (anonymous):

i got 45

OpenStudy (haseeb96):

yes u are correct let me review it again

OpenStudy (haseeb96):

(f*g)(x) = f(g(x)) = f(2x-3) = 4(2x-3)-7 = 8x -12 -7 = 8x - 19 now replace x by 4 in (f*g)(4) = 8(4) -19 = 32-19 = 13 is correct answer

OpenStudy (haseeb96):

Sorry for misunderstanding I was busy in chatting with my friends sorry again

OpenStudy (haseeb96):

@Emilys819 Did u get it ? now

OpenStudy (anonymous):

no prob i see exactly what i need to do now thankyou!

OpenStudy (anonymous):

yes i did

OpenStudy (haseeb96):

Thank you. Sorry again

OpenStudy (anonymous):

completely fine